OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL!!! the sum of two numbers is 23 and the product of two numbers is 60 what are the two numbers?

OpenStudy (mathstudent55):

Let the numbers be x and y. How do you write the sum of x and y is 23?

OpenStudy (anonymous):

x+y=23?

OpenStudy (anonymous):

i have to finish in 3 minutes

OpenStudy (anonymous):

I can't work in a rush Sorry

OpenStudy (mathstudent55):

Yes. Now you need the product of the numbers is 60: xy = 60

OpenStudy (mathstudent55):

You have a system of equations: x + y = 23 xy = 60

OpenStudy (mathstudent55):

Solve first eq for x: x = y - 23 Replace into x of second eq.: (y - 23)y = 60 y^2 - 23y - 60 = 0